# StyleReiser: Stylizing Video With Reinforced Structure Guide

### v1

@@ Abstract
We introduce StyleReiser, an example-based video stylization method that transfers style from a given keyframe to the entire video sequence while maintaining visual consistency even in distant frames where the scene structure may change significantly. Unlike previous keyframe-based methods, our approach considers consistency with the prescribed style and maintains fidelity to new structural elements appearing in the target video sequence. This combination can significantly improve the quality of the stylized sequence without the need to add more correction keyframes. We also demonstrate that our approach can notably enhance the output of text-driven video stylization methods by suppressing their structural instability and enabling the user to perform custom edits on the generated keyframes. Moreover, due to its capability to perform inference in real-time, our technique can also be applied in interactive scenarios, such as consistently stylized video calls, which are difficult to achieve using text-driven approaches.
@@ 1 Introduction
Conditional video generation has received considerable attention recently. Given a combination of content and structure guidance, approaches that utilize large text-to-image diffusion models have achieved impressive results Yang et al. (2023); Ceylan et al. (2023); Geyer et al. (2024). However, the present video-to-video frameworks struggle with two significant issues: (i) poor temporal consistency of the stylized output, and (ii) insufficient fidelity to the input video (cf. Fig. 1d). These shortcomings can significantly impact the quality and usability of the generated content. Furthermore, to generate a single frame, considerable computational resources are required as a result of the use of diffusion models. Another important drawback of diffusion models is the limited control over stylized content. It is often difficult to come up with a text prompt that expresses exactly the intended artistic vision. In some cases manual intervention is inevitable, and it is not clear how to incorporate those additional edits. On the other hand, feed-forward video-to-video methods that are not dependent on diffusion models Jamriška et al. (2019); Texler et al. (2020b); Futschik et al. (2021) excel in semantically meaningful stylization, effectively preserving key features of the exemplar style. Users are granted full creative freedom with the ability to directly manipulate the style exemplar, bypassing the need for laborious refinement of text prompts, and have interactive feedback Texler et al. (2020b). However, when a new structural element appears in the input video, their results degrade considerably (cf. Fig. 1e), and additional stylized keyframes need to be provided. Those can be tedious to create manually or can introduce inconsistencies when generative methods are used for their preparation.
In this paper, we demonstrate that a key limitation of previous keyframe-based video stylization approaches lies in the fact that they are aiming to transfer the style but not explicitly trying to preserve structure of the target content. This omission may cause difficulties when significant changes occur in the frames farther away from the keyframe. To mitigate this drawback, we propose to relax on style guidance and focus more on the increased fidelity to novel structural elements appearing in the input video. This helps to significantly prolong the resistance to structural changes (cf. Fig. 1c) and thus eliminate the need to specify additional keyframes. Such an improvement brings benefits for both manually prepared keyframes where it saves time and effort of an artist as well as for synthetically generated ones where it prevents unwanted injection of structural inconsistencies caused by diffusion models. In addition, since our trained network can process the input video in real time, we can significantly reduce computational overhead in cases such as a video conference scenario (as was already demonstrated, e.g., in Futschik et al. Futschik et al. (2021)) where diffusion-based methods are difficult to apply.
@@
Figure 1: An example of style transfer with reinforced structure guidance. We train a feed-forward neural network to transfer the style of a single stylized keyframe (a) to the remaining video frames (b) producing the output stylized frames (c). The keyframe (a) was created using text-driven video-to-video diffusion model of Ceylan et al. Ceylan et al. (2023). When the same model is applied on a different frame from the same video sequence (b), a significant inconsistency in structure can occur (d). The same holds for the result of state-of-the-art keyframe-based video stylization method Futschik et al. (2021). In contrast our approach (c) offers a stylized result that is faithful to the structural changes in the target video sequence and thus can deliver consistent output.
@@ 2 Related work
Figure 2: An overview of our approach. Given images from a source domain $\mathbf{x}_{i}\in\mathcal{X}$, we optimize the network $f$ to produce images $\hat{\mathbf{y}}_{i}$ with a similar appearance as images from a target domain $\mathbf{y}_{k}$. The key loss $\mathcal{L}_{\text{key}}$ (1) encourages reconstruction of keyframes $\mathbf{y}_{k}$, the VGG loss $\mathcal{L}_{\text{VGG}}$ (2) ensures style consistency between frames and keyframes using Gram correlation matrices $g$ of extracted VGG network responses $v$ Gatys et al. (2016), and finally the cSDS loss $\mathcal{L}_{\text{cSDS}}$ (3) forces fidelity to structural elements present in the input video frames $\mathbf{x}_{i}$. cSDS loss requires a pre-trained ControlNet Zhang et al. (2023b) model consisting of diffusion model $d$ initialized by adding a random Gaussian noise $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ into the synthesized image $\hat{\mathbf{y}}_{i}$, a time step $t$, and a function $c$ transforming the input image $\mathbf{x}_{i}$ to a condition $\mathbf{c}$ (in this case Canny edge detector Canny (1986)).
@@
3 Method
A basic overview of our approach is given in Fig. 2. We train a consistent stylization operator (a modified variant of U-Net Ronneberger et al. (2015)) to stylize all frames of an input video sequence, given one (or possibly more) stylized keyframes. Those keyframes can be prepared manually, through a text-to-image generative model, or using a combination of both approaches, i.e., generating a stylized version of an input keyframe and then correcting it manually.
For the stylization operator $f$, we adopt the network architecture design proposed by Futschik et al. Futschik et al. (2019) that proved to be suitable for style transfer tasks Texler et al. (2020a); Futschik et al. (2021) allowing for reproduction of important high-frequency details, critical for generating complex and believable artistic styles.
$\mathbf{x}$
@@ 3 Method
Figure 3: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization: The target video sequence (see a representative target frame $\mathbf{x}$) has been stylized using diffusion-based approaches (top row): (a) Ceylan et al. Ceylan et al. (2023), (b) Yang et al. Yang et al. (2023), (c) Chu et al. Chu et al. (2024), and (d) Geyer et al. Geyer et al. (2024). One frame from those stylized sequences was used as a keyframe (see small insets). The style of this keyframe has been propagated to the rest of the target sequence $\textbf{x}\in\mathcal{X}$ using our approach (bottom row). Note how our approach better preserves the structural details seen in the target frame. Also, see our supplementary video to compare consistency across the entire sequence. Diffusion-based approaches tend to suffer from notable structural flicker, whereas our approach keeps the structure consistent, yielding considerably more stable results.
Given two sets of images, we train $f$ to perform a consistent stylization on both sets. The first set $\mathcal{K}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathcal{X}\times\mathcal{Y}\mid i\in\mathbf{K}\}$ of keyframes contains $K$ tuples of original frames $\mathbf{x}\in\mathcal{X}$ and their stylized counterparts $\mathbf{y}\in\mathcal{Y}$. The symbol $\mathcal{X}$, $|\mathcal{X}|=N$, denotes a set of all input images, the symbol $\mathcal{Y}$, $|\mathcal{Y}|=K$, denotes a set of all stylized images, and the symbol $\mathbf{K}$, $|\mathbf{K}|=K$, denotes a set of frame indices to which a ground truth stylized version exists. The second set $\mathcal{Z}=\{\mathbf{x}_{i}\in\mathcal{X}\mid i\not\in\mathbf{K}\}$, $|\mathcal{Z}|=N-K$, consists of all frames of the input video to which there is no stylized counterpart. Both sets of images are used during the training step of our approach.
Let $f(\mathbf{x}_{i};\mathbf{\Phi})$ be the output of our neural network for the $i$-th input image and $\mathbf{\Phi}$ all the learnable parameters of the network. We use an $\ell_{2}$ loss
@@
$\mathcal{L}_{\text{key}}(\mathcal{K};\mathbf{\Phi})=\frac{1}{|\mathbf{K}|}\sum_{i\in\mathbf{K}}\left\|f\left(\mathbf{x}_{i};\mathbf{\Phi}\right)-\mathbf{y}_{i}\right\|_{2}^{2}$ (1)
to encourage the $f$ operator towards reconstruction of the user-provided stylized keyframes.
In addition to the $\ell_{2}$ loss, we apply a distribution matching, VGG-based, style loss, computed as
$\begin{split}\mathcal{L}_{\text{VGG}}(\mathcal{K},\mathcal{Z};\mathbf{\Phi},\mathbf{\Psi})&=\\ =\frac{1}{N|L|}\sum_{i\not\in\mathbf{K}}\sum_{j\in\mathbf{K}}&\sum_{l\in\mathbf{L}}\left\|g(v(f(\mathbf{x}_{i};\mathbf{\Phi}),l;\mathbf{\Psi}))-g(v(\mathbf{y}_{j},l;\mathbf{\Psi}))\right\|_{2}^{2}\end{split}$ (2)
where $g(\mathbf{\cdot})$ computes the Gram correlation matrix Gatys et al. (2016), $v(\mathbf{\cdot},l;\mathbf{\Psi})$ extracts VGG network responses Simonyan and Zisserman (2014) at the layer $l$ with the VGG network parameters $\mathbf{\Psi}$, and $\mathbf{L}$ is a set of indices of all considered VGG layers. Distribution matching by minimizing differences between Gram matrices of neural VGG feature responses was originally proposed by Gatys et al. Gatys et al. (2016) for an optimization-based image stylization framework. Exactly as in their optimization framework, the parameters of the VGG network $\mathbf{\Psi}$ are frozen during the training of our operator $f$.
The last term of our operator loss is a variant of Score Distillation Sampling (SDS) proposed by Poole et al. Poole et al. (2022). SDS utilizes an off-the-shelf pre-trained text-to-image diffusion model to serve as a regularization technique in pursuit of realistically looking images. An image diffusion model is a generative model trained to invert the diffusion process, which progressively corrupts a natural image into images of statistically perfect Gaussian noise Ho et al. (2020). By removing small amounts of Gaussian noise with a fixed variance schedule $\bar{\alpha}_{t}$, an image is denoised. In their original work Poole et al. (2022), given a randomly sampled time step $t\sim\mathcal{U}(0,1)$, a random Gaussian noise $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, a noisy version of an image $\hat{\mathbf{y}}_{i}=f(\mathbf{x}_{i};\mathbf{\Phi})$, $\hat{\mathbf{y}}_{i,t}=\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{y}}_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon}$, and textual description $\mathbf{p}$, they show that a pre-trained text-to-image diffusion model $d(\hat{\mathbf{y}}_{i,t},\mathbf{p},t;\mathbf{\Theta)}$, parameterized by $\mathbf{\Theta}$, can be successfully utilized for optimization problems. In contrast to their work, we executed two key modifications of SDS. First, we made $t$ a constant. As shown in our experiments, the value of $t$ contributes to setting the level of fidelity to the input video structures. Second, we replaced the textual guidance by an image guidance. We set textual description $\mathbf{p}$ to an empty string and replaced a standard text-to-image diffusion model with a ControlNet Zhang et al. (2023b) additionally conditioned on $\mathbf{c}$ resulting in controlled Score Distillation Sampling (cSDS)
$\mathcal{L}_{\text{cSDS}}(\mathcal{Z};\mathbf{\Phi},\mathbf{\Theta})=\frac{1}{N}\sum_{i\not\in\mathbf{K}}\left\|d\left(\hat{\mathbf{y}}_{i,t},\mathbf{p},\mathbf{c},t;\Theta\right)-\epsilon\right\|_{2}^{2}.$ (3)
Note that the parameters $\mathbf{\Theta}$ of the “controlled” denoising diffusion model itself are fixed during the training of our operator. Also note that in practice, the back-propagation through the diffusion model is skipped to speed up the training. The condition $\mathbf{c}$ is the input image $\mathbf{x}$ transformed by a predefined function $c(\mathbf{x})$, e.g., Canny edge detector Canny (1986). The total loss for training the operator becomes
$\begin{split}\mathcal{L}(\mathcal{K},\mathcal{Z};\mathbf{\Phi},\mathbf{\Psi})=&\lambda_{k}\mathcal{L}_{\text{key}}(\mathcal{K};\mathbf{\Phi})+\\ +&\lambda_{v}\mathcal{L}_{\text{VGG}}(\mathcal{K},\mathcal{Z};\mathbf{\Phi},\mathbf{\Psi})+\\ +&\lambda_{c}\mathcal{L}_{\text{cSDS}}(\mathcal{Z};\mathbf{\Phi},\mathbf{\Theta}).\end{split}$ (4)
$\mathbf{x}$
@@
keyframe from Yang et al. Yang et al. (2023) $(\mathbf{x}_{r},\mathbf{y}_{r})$
custom edit $\mathbf{y}_{e}$
target frame $\mathbf{x}$
@@
Figure 7: Custom edit of results generated by the diffusion-based method of Yang et al. Yang et al. (2023). Left to right: (a, b) keyframe $(\mathbf{x}_{r},\mathbf{y}_{r})$$\mathbf{y}_{r}$ is the result of Yang et al. Yang et al. (2023) conditioned with the text prompt “Galadriel, the royal Elf, silver-golden hair,” (c) custom edit $\mathbf{y}_{e}$ of the stylized keyframe $\mathbf{y}_{r}$, (d) target frame $\mathbf{x}$, (e) stylization $\hat{\mathbf{y}}$ produced by our method with a single keyframe $(\mathbf{x}_{r},\mathbf{y}_{e})$. A key advantage of our method is that it allows custom edits of videos stylized by diffusion-based methods, which provide only a limited textual control over the generated results.
target frame $\mathbf{x}$
@@
target frame $\mathbf{x}$
@@
target frame $\mathbf{x}$

### v2

# StructuReiser: A Structure-preserving Video Stylization Method

@@ Abstract
We introduce StructuReiser, a novel video-to-video translation method that transforms input videos into stylized sequences using a set of user-provided keyframes. Unlike existing approaches, StructuReiser maintains strict adherence to the structural elements of the target video, preserving the original identity while seamlessly applying the desired stylistic transformations. This enables a level of control and consistency that was previously unattainable with traditional text-driven or keyframe-based methods. Furthermore, StructuReiser supports real-time inference and custom keyframe editing, making it ideal for interactive applications and expanding the possibilities for creative expression and video manipulation.
@@ 1 Introduction
Guided video stylization has received considerable attention recently. The example-based methods Jamriška et al. (2019); Texler et al. (2020b); Futschik et al. (2021) offer users full creative freedom by allowing them to directly manipulate the appearance through one or more stylized keyframes. Alternatively, diffusion models Yang et al. (2023); Ceylan et al. (2023); Geyer et al. (2024) provide a simplified workflow, enabling users to modify appearance using various text prompts.
Despite their practicality and impressive results, both keyframe-based and text-driven approaches share a significant limitation that can impact the quality and usability of generated content: they transfer the style without adequately preserving the structure of the target content. For example, when the input video features a distinct character whose identity is crucial (see Fig. 1a), there is no guarantee that the stylized sequence will retain this identity (see Fig. 1d–e). To address this issue in keyframe-based approaches, users must provide a set of keyframes that accurately capture existing and newly appearing structural elements in the input video – a process that can be tedious and time-consuming. The challenge is even greater with text-driven methods, where crafting a prompt that ensures the preservation of structural details is nearly impossible.
To overcome these challenges, we propose a novel approach to keyframe-based video stylization. The core idea of our solution is a new formulation of guided video stylization that, in addition to ensuring fidelity to the transferred style, also focuses on preserving the structural elements of the input video (cf. Fig. 1b–c). To the best of our knowledge, this is the first attempt to address structural fidelity in both text-driven and keyframe-based video stylization. This improvement not only reduces the manual effort required for keyframe preparation, but also prevents the unintended introduction of structural inconsistencies that can arise in text-driven approaches.
To validate the effectiveness of our method, we conducted extensive qualitative and quantitative evaluations, including an online user survey, which demonstrated significant improvements over the current state-of-the-art. Moreover, our method supports real-time inference, making it suitable for interactive scenarios such as video conferencing, where text-driven approaches are difficult to apply.
To summarize our contributions:
  1. (i)
    We formulate the task of structurally faithful stylization of videos, establishing a practical framework that emphasizes both stylistic fidelity and structural preservation.
  2. (ii)
    Building on this framework, we introduce a novel keyframe-based video stylization method that effectively transfers style from a single stylized keyframe while preserving essential structural elements in the entire video sequence.
  3. (iii)
    We validate our approach through extensive qualitative and quantitative evaluations, showing a significant improvement over current state-of-the-art in maintaining structural fidelity while closely adhering to the desired stylization.
@@
Figure 1: StructuReiser transfers the style from a single stylized keyframe (a) to the entire video sequence (b) generating stylized frames (c) that are both stylistically consistent and structurally faithful. The keyframe (a) was created using the text-guided video-to-video diffusion model by Ceylan et al. Ceylan et al. (2023). However, when applied directly to other frames in the sequence, this model often introduces significant structural inconsistencies (d). The state-of-the-art keyframe-based video stylization method of Futschik et al. Futschik et al. (2021) faces similar issues (e). In contrast, our approach (c) maintains the structural integrity of the target video sequence while ensuring coherent stylization throughout.
@@ 2 Related work
Figure 2: An overview of our approach. Given images from a source domain $\mathbf{y}_{i}\in\mathcal{Y}$, we optimize the operator $f$ to produce images $\hat{\mathbf{y}}_{i}$ with a similar appearance as images from a target domain $\hat{\mathbf{x}}_{i}$. The key loss $\mathcal{L}_{\text{key}}$ (2) encourages reconstruction of keyframes $\hat{\mathbf{x}}_{i}$, the style loss $\mathcal{L}_{\text{style}}$ (3) ensures style consistency between frames and keyframes using Gram correlation matrices $g$ of extracted VGG network responses $v$ Gatys et al. (2016), and finally the structure loss $\mathcal{L}_{\text{structure}}$ (4) enforces fidelity to structural elements present in the input video frames $\mathbf{y}_{i}$. The structure loss requires a pre-trained ControlNet Zhang et al. (2023b) model consisting of diffusion model $d$ initialized by adding a random Gaussian noise $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ into the synthesized image $\hat{\mathbf{y}}_{i}$, a time step $t$, and a function $c$ transforming the input image $\mathbf{y}_{i}$ to a condition $\mathbf{c}$ (in this case Canny edge detector Canny (1986)).
@@
3 Our Approach
The input of our method is a set of $N$ video frames $\mathcal{T}$ which is partitioned into two subsets: $\mathcal{T}=\mathcal{X}\cup\mathcal{Y}$.
  1. 1.
    Keyframes $\mathcal{X}=\{\mathbf{x}_{i}\in\mathcal{T}\mid i\in\mathbf{K}\}$, where $\mathbf{K}$ is a set of $K$ frame indices. For each keyframe $\mathbf{x}_{i}$, a ground truth stylized version  $\mathbf{\hat{x}}\in\mathcal{\hat{X}}$ is provided.
  2. 2.
    Non-keyframes $\mathcal{Y}=\{\mathbf{y}_{i}\in\mathcal{T}\mid i\not\in\mathbf{K}\}$, consisting of the remaining $N-K$ frames without stylized counterparts.
We define the set of keyframe pairs as $\mathcal{K}=\{(\mathbf{x}_{i},\mathbf{\hat{x}}_{i})\in\mathcal{X}\times\mathcal{\hat{X}}\mid i\in\mathbf{K}\}$, which contains $K$ tuples of original keyframes and their corresponding stylized versions.
The aim of our approach is to learn a stylization operator $f$ that maps any input video frame $\mathbf{z}_{i}\in\mathcal{Z}$ into its stylized counterpart $\hat{\mathbf{z}}_{i}\in\mathcal{\hat{Z}}$, i.e., $\hat{\mathbf{z}}_{i}=f(\mathbf{z}_{i})$, by minimizing the following loss function:
$\mathcal{L}(\mathcal{K},\mathcal{Y})=\lambda_{k}\mathcal{L}_{\text{key}}(\mathcal{K})+\lambda_{v}\mathcal{L}_{\text{style}}(\mathcal{K},\mathcal{Y})+\lambda_{s}\mathcal{L}_{\text{structure}}(\mathcal{Y}).$ (1)
$\mathbf{y}$
@@ 3 Our Approach
Figure 3: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization: The target video sequence (see a representative target frame $\mathbf{y}$) has been stylized using diffusion-based approaches (top row): (a) Ceylan et al. Ceylan et al. (2023), (b) Yang et al. Yang et al. (2023), (c) Chu et al. Chu et al. (2024), and (d) Geyer et al. Geyer et al. (2024). One frame from those stylized sequences was used as a keyframe (see small insets). The style of this keyframe has been propagated to the rest of the target sequence $\textbf{y}\in\mathcal{Y}$ using our approach (bottom row). Note how our approach better preserves the structural details seen in the target frame. Also, see our supplementary video to compare consistency across the entire sequence. Diffusion-based approaches tend to suffer from notable structural flicker, whereas our approach keeps the structure consistent, yielding considerably more stable results.
The reconstruction loss $\mathcal{L}_{\text{key}}$ guides the $f$ operator to generate outputs that closely match the user-provided stylized keyframes, i.e.:
$\mathcal{L}_{\text{key}}(\mathcal{K})=\frac{1}{K}\sum_{i\in\mathbf{K}}\left\|f\left(\mathbf{x}_{i}\right)-\mathbf{\hat{x}}_{i}\right\|_{2}^{2}.$ (2)
The style loss $\mathcal{L}_{\text{style}}$ originally proposed by Gatys et al. Gatys et al. (2016) ensures that the distribution of stylized frames $f(\mathbf{y}_{i})\in\mathcal{\hat{Y}}$ aligns with that of the stylized keyframes $f(\mathbf{x}_{i})\in\mathcal{\hat{X}}$. This is achieved by comparing responses of the VGG network Simonyan and Zisserman (2014) as follows:
$\mathcal{L}_{\text{style}}(\mathcal{K},\mathcal{Y})=\frac{1}{|\mathbf{K}||\mathbf{L}||\mathcal{Y}|}\sum_{i\in\mathbf{K}}\sum_{j\not\in\mathbf{K}}$ $\sum_{l\in\mathbf{L}}\left\|g\left(\mathbf{\hat{x}}_{i},l\right)-g\left(f(\mathbf{y}_{j}),l\right)\right\|_{2}^{2}$ (3)
where $g(\mathbf{\cdot})$ computes the Gram correlation matrix of VGG network responses at the layer $l\in\mathbf{L}$ and $\mathbf{L}$ is a set of indices of all considered VGG layers.
The structure loss $\mathcal{L}_{\text{structure}}$ is the key component and the main contribution of our framework. It guides the operator $f$ in transferring essential structural elements from the input video sequence to the stylized output. We compute this loss using a diffusion model $d$, conditioned by $c$, to ensure preservation of important structural details:
@@
$\mathcal{L}_{\text{structure}}(\mathcal{Y})=\frac{1}{|\mathcal{Y}|}\sum_{i\not\in\mathbf{K}}\left\|d\left(\hat{\mathbf{y}}_{i,t},c(\textbf{y}_{i}),t\right)-\epsilon\right\|_{2}^{2},$ (4)
where $\hat{\mathbf{y}}_{i,t}=\sqrt{\bar{\alpha}_{t}}f(\mathbf{y}_{i})+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon}$ is a noisy version of the stylized input frame $\mathbf{y}_{i}$ corrupted by a Gaussian noise $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at a pre-defined time step $t$.
In Fig. 1c, $d$ is a pre-trained diffusion model ControlNet Zhang et al. (2023b), conditioned by a line art detector $c$. However, this is just one of many possible choices. We further explore different conditionings $c$ of the diffusion model $d$ in Sec. 5.4.
Our strategy to computing $\mathcal{L}_{\text{structure}}$ shares some aspects with the score distillation sampling (SDS) method proposed by Poole et al. Poole et al. (2022), however, a key difference in our solution is its ability to preserve prescribed structural details, which is not possible with the original SDS. In Sec. 5.5 we also demonstrate the importance of this particular formulation in contrast to a straightforward approach, i.e., directly computing the structure loss using a line art detector.
Definition of “structure”
We use terms like “structure” and “structural fidelity” to describe key visual elements from the input video we aim to preserve. However, we acknowledge that precise definitions of these terms are challenging and interpretations may vary. Fortunately, “structure” is defined inherently in our approach – by the choice of conditioning $c$ in the diffusion model $d$, determining the meaning of “structure” by design.
SSIM $\uparrow$ LPIPS $\downarrow$
F
LIP $\downarrow$
ours 0.75 $\pm$ 0.12 0.28 $\pm$ 0.09 0.30 $\pm$ 0.06
Geyer et al. Geyer et al. (2024) 0.72 $\pm$ 0.10 0.30 $\pm$ 0.08 0.31 $\pm$ 0.05
ours 0.71 $\pm$ 0.10 0.37 $\pm$ 0.09 0.43 $\pm$ 0.08
Yang et al. Yang et al. (2023) 0.65 $\pm$ 0.09 0.42 $\pm$ 0.09 0.47 $\pm$ 0.08
ours 0.71 $\pm$ 0.09 0.38 $\pm$ 0.10 0.40 $\pm$ 0.10
Ceylan et al. Ceylan et al. (2023) 0.67 $\pm$ 0.08 0.41 $\pm$ 0.10 0.42 $\pm$ 0.10
ours 0.62 $\pm$ 0.07 0.48 $\pm$ 0.10 0.59 $\pm$ 0.11
Chu et al. Chu et al. (2024) 0.56 $\pm$ 0.07 0.55 $\pm$ 0.07 0.61 $\pm$ 0.10
ours 0.69 $\pm$ 0.14 0.33 $\pm$ 0.12 0.37 $\pm$ 0.08
Jamriška et al. Jamriška et al. (2019) 0.67 $\pm$ 0.11 0.36 $\pm$ 0.10 0.38 $\pm$ 0.06
Futschik et al. Futschik et al. (2021) 0.66 $\pm$ 0.11 0.35 $\pm$ 0.11 0.39 $\pm$ 0.07
Texler et al. Texler et al. (2020b) 0.66 $\pm$ 0.11 0.36 $\pm$ 0.09 0.38 $\pm$ 0.06
Table 1: Quantitative comparison of structural fidelity. Averages and standard deviations between input and stylized frames shown in five groups. Since diffusion-based methods cannot perform exemplar-based stylization, we trained our method using style exemplars generated by these methods for the first four groups (see Sec. 4.2). The fifth group includes exemplar-based approaches trained on a union of the diffusion-generated and artist-created exemplars. Our method outperforms existing approaches across all metrics.
$\mathbf{y}$
@@
keyframe from Yang et al. Yang et al. (2023) $(\mathbf{x}_{r},\hat{\mathbf{x}}_{r})$
custom edit $\hat{\mathbf{x}}_{e}$
target frame $\mathbf{y}$
@@
Figure 7: Custom edit of results generated by the diffusion-based method of Yang et al. Yang et al. (2023). Left to right: (a, b) keyframe $(\mathbf{x}_{r},\hat{\mathbf{x}}_{r})$ is the result of Yang et al. Yang et al. (2023) conditioned with the text prompt “Galadriel, the royal Elf, silver-golden hair,” (c) custom edit $\hat{\mathbf{x}}_{e}$ of the stylized keyframe $\hat{\mathbf{x}}_{r}$, (d) target frame $\mathbf{y}$, (e) stylization $\hat{\mathbf{y}}$ produced by our method with a single keyframe $(\mathbf{x}_{r},\hat{\mathbf{x}}_{e})$. A key advantage of our method is that it allows custom edits of videos stylized by diffusion-based methods, which typically offer only limited control over the generated results through textual prompts.
target frame $\mathbf{y}$
@@
target frame $\mathbf{y}$
@@
target frame $\mathbf{y}$
@@
3.1 Implementation details
We implemented our approach in PyTorch Paszke et al. (2019) using the AdamW Loshchilov and Hutter (2017) optimizer with fixed learning rate $3\cdot 10^{-5}$. To model the stylization operator $f$, we adopt the network architecture originally proposed by Futschik et al. Futschik et al. (2019) that proved to be suitable for style transfer tasks Texler et al. (2020a); Futschik et al. (2021) allowing for reproduction of important high-frequency details, critical for generating complex and believable artistic styles. The batch size was set to $1$ and instance normalization Ulyanov et al. (2016) used instead of batch normalization. We set $\lambda_{k}=1.0$ and $\lambda_{v}=100.0$. The parameters for $\mathcal{L}_{\text{structure}}$ (4) were selected experimentally and differ between the presented sequences: $\lambda_{s}\in\{10^{-5},10^{-6}\}$ and $t\in\{20,28\}$. As $d$, we adopt the default noise scheduler of ControlNet v1.1 Zhang et al. (2023b) and the UniPC scheduler Zhao et al. (2024) with the number of steps set to $30$. The parameters of the VGG network and ControlNet were fixed and line art conditioning was used for $c$ (if not stated otherwise). The training was performed on a single NVIDIA A100 GPU with $40$ GB of RAM for $4$ hours and the model with the lowest total loss has been selected to produce the stylized sequences.
0.0
0.2
0.4
0.6
0.8
1.0
Temporal consistency $\square$
0.0
0.2
0.4
0.6
0.8
1.0
Style
Ours vs. Geyer et al. Geyer et al. (2024)
Ours vs. Yang et al. Yang et al. (2023)
Ours vs. Chu et al. Chu et al. (2024)
Ours vs. Ceylan et al. Ceylan et al. (2023)
Ours vs. Futschik et al. Futschik et al. (2021)
Ours vs. Texler et al. Texler et al. (2020b)
Ours vs. Jamriška et al. Jamriška et al. (2019)
0.0
0.2
0.4
0.6
0.8
1.0
Structure $\triangle$
Figure 11: Perceptual study. Each point represents the ratio of votes preferring the results of our method over those of other methods, based on responses from a total of $54$ participants. Comparisons were made against four diffusion-based methods – Geyer et al. Geyer et al. (2024) (pink), Yang et al. Yang et al. (2023) (orange), Chu et al. Chu et al. (2024) (brown), Ceylan et al. Ceylan et al. (2023) (purple) – and three exemplar-based methods – Futschik et al. Futschik et al. (2021) (green), Texler et al. Texler et al. (2020b) (blue), and Jamriška et al. Jamriška et al. (2019) (red). The bottom x-axis displays the ratio of answers favoring our method for preserving temporal consistency (hollow squares), the top x-axis shows the ratio favoring our method for structure preservation (hollow triangles), and the y-axis represents the style reproduction ratio. The graph illustrates that our method significantly outperforms previous works in reproducing input structures and maintaining temporal consistency. Conversely, prior methods are preferred for style preservation, which is expected given our focus on structural fidelity.